A particle is projected with a velocity u at an angle alpha with the horizontal. At what time its vertical component of velocity is half of its net speed at heighest point
 

Grade:11

1 Answers

Khimraj
5 years ago
net speed at highest  point = ucos\alpha
then
v = u -gt
ucos\alpha/2 = usin\alpha – gt
Then
t = usin\alpha/g – ucos\alpha/2g
Hope it clears.